OpenStudy (anonymous):

The tangent line to y= f(x) at (4,-6) passes through the point (10,9) Compute the following, f(4) and f'(4)

OpenStudy (anonymous):

well you know that the slope of the tangent line at the point (4,-6) is f'(4) ?

OpenStudy (anonymous):

Right!

OpenStudy (anonymous):

so we can compute f'(4) using the two points the tangent line passes through : (4,-6) ,(10,9) can you do it ?

OpenStudy (anonymous):

Maybe... do I just find the slope using the y-y/x-x formula and thats my answer?

OpenStudy (anonymous):

yes it is ! (for f'(4))

OpenStudy (anonymous):

awesome! now I'm completely lost on f(4) haha

OpenStudy (anonymous):

well .. f(4) is right in front of your eyes! you got the point with x=4 that f passes through it!!

OpenStudy (anonymous):

okay I'm completely confused.. I'm sorry! I'm way too new at this haha. I just do y=mx+b and plug in 4 for x, the answer to f'(4) for the slope?

OpenStudy (anonymous):

f'(4) is the slope of the tangent line.. we know that the tangent line passes through the points : (4,-6) ,(10,9) so .. all you have to do is calculate the slope of the line that passes through those two points and get f'(4) (just as you said up there)

OpenStudy (anonymous):

f'(4) = (9 - (-6)) / (10 - 4) = 15 / 6 = 2.5

OpenStudy (anonymous):

you DONT need to find the line equation y=mx+b for this question.. you just need the slope of the tangent line (f'(4))

OpenStudy (anonymous):

Alright, I think I get it! Thanks a bunch!

OpenStudy (anonymous):

wait.. do you know what is f(4) ?

OpenStudy (anonymous):

it's easier than f'(4)

OpenStudy (anonymous):

it's given in the question!

OpenStudy (anonymous):

-6?

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

yes.. very good :)